# Rings of Hilbert modular forms, computations on Hilbert modular surfaces, and the Oda-Hamahata conjecture

### v1

@@ 1. Introduction
Let $E$ be an elliptic curve over a real quadratic field belonging to an isogeny class with one of the following LMFDB labels or their Galois conjugates: 2.2.5.1/31.1-a, 2.2.5.1/55.1-a, 2.2.5.1/80.1-a, 2.2.8.1/9.1-a, 2.2.8.1/17.1-a, 2.2.8.1/32.1-a, 2.2.13.1/9.1-a, 2.2.17.1/4.1-a, 2.2.17.1/8.3-a. Then there is a correspondence (not necessarily defined over ${\mathbb{Q}}$) between the Hilbert modular variety of level equal to the conductor of $E$ and the Kummer surface of $E\times E^{\sigma}$.
@@
We now return to elliptic curves over $K$. Let $E$ be such a curve and let its conductor be $I$. To generalize the algebraic statement of modularity is fairly straightforward: one requires that the coefficients of the $L$-function of $E$ match the Hecke eigenvalues of a Hilbert modular form $f$. This is known for all elliptic curves over real quadratic fields and many over fields of higher degree: see [10, Theorems 1-2].
@@
[14, Section 0, p. 194] Let $E$ be an elliptic curve over a real quadratic field $K$ with conductor $I$. Then there is a correspondence between $H_{I,K}$ and $K_{E}=(E\times E^{\sigma})/\pm 1$ that induces an injection from the primitive part of $\mathcal{H}^{2}(K_{E})$ to that of $\mathcal{H}^{2}(H_{I,K})$.
@@
Hamahata’s conjecture is known under some special circumstances: in particular, if $E$ is a ${\mathbb{Q}}$-curve, i.e., an elliptic curve isogenous to all of its Galois conjugates. This is not clearly stated or proved in the literature, but is known to experts. It is a consequence of the modularity of $E$, which is proved in [27, Corollary 6.2], subject to Serre’s conjecture on modular $2\times 2$ Galois representations, which is now a theorem of Khare and Wintenberger [17] More precisely, Ribet’s result is that $E$ is a quotient of $J_{1}(N)$; it follows that there is a dominant map $J_{1}(N)^{d}\to\prod_{i=1}^{d}E^{\sigma_{i}}$, where $d=[K:{\mathbb{Q}}]$. The domain of this map is related to $\operatorname{H}^{d}(X_{0}(N))$ by Hecke correspondences, and the equality of the levels follows from Serre’s formula for the level of the modular form, also proved by Khare and Wintenberger.
@@
Let ${\mathcal{O}}$ be an order in a totally real field and ${\mathcal{O}}^{+}$ the set of its totally nonnegative elements. Let ${\mathfrak{d}}_{\mathcal{O}}$ be the different of ${\mathcal{O}}$. A Hilbert modular form for ${\mathcal{O}}$ has a Fourier expansion in which the coefficients are indexed by elements of ${\mathcal{O}}^{+}\cap{\mathfrak{d}}_{\mathcal{O}}^{-1}$ and are invariant under multiplication by elements of ${\mathcal{O}}^{+}\cap{\mathcal{O}}^{*}$. In order to multiply two such expansions, one needs an algorithm for writing an element of ${\mathcal{O}}^{+}$ as a sum of two such elements in all possible ways. An efficient algorithm for doing so will be described in [1].
@@
In particular, this theorem applies in all cases where $H_{I,K}$ is not of general type and $I\neq(1)$, except for $K={\mathbb{Q}}(\sqrt{13}),I=(2)$. In this case we do find a K3 surface that is in correspondence with the Hilbert modular variety (and in fact this has been known for some years, being [35, Theorem 9]) but it is quite difficult to establish an explicit correspondence with the desired Kummer surface. In two of the cases where $I=(1)$ the birational equivalence between the Hilbert modular surface and a K3 surface with canonical singularities is made explicit in [37, Appendix].
@@
Our method is to study the Baily-Borel compactification $H_{I,K}$ by means of computer algebra and exhibit the desired correspondence to the Kummer surface by explicit maps. In some cases the Kummer surface of $E\times E^{\sigma}$ is itself a quotient of $H_{I,K}$; in other cases this does not appear to be true and instead we find an auxiliary K3 surface $S$ that admits maps of finite degree from both $H_{I,K}$ and the Kummer surface of $E\times E^{\sigma}$. The methods are similar to those of [37, Appendix], except that we do not necessarily start with a variety birational to a K3 surface. In addition, the setup of [37] requires information about the ring of Hilbert modular forms that is not available here: in particular, that it can be generated by Borcherds products. Against this, we do not find equations for the Hirzebruch-Zagier cycles [36, Def. VI.1.4] on the Hilbert modular surfaces.
@@
In other cases, the surface or a quotient has geometric genus $1$ but is of general type. It is known [23, Theorem 1] that if a surface $S$ is of geometric genus $1$ then there is a K3 surface $K_{S}$ and a correspondence between $S$ and $K_{S}$ identifying the transcendental part of $H^{2}_{{\mathrm{\acute{e}t}}}$ for the two surfaces. However, finding such a K3 surface explicitly is a difficult and interesting problem in general. In some such cases the quotient by some or all of the Atkin-Lehner involutions is a surface with $p_{g}=1$ and $\kappa<2$ and we can find the desired correspondence, but otherwise there is no obvious method to use. This problem will be studied in future work.
@@ 2.2. Projective embedding and cusps
It can happen that we run out of room, in the sense that projection away from a singular point or line gives a map of degree greater than $1$. This is undesirable, because the image may no longer have the desired cohomology. When it happens, we apply the $2$-fold Veronese embedding and proceed with our projections as before.
@@ 2.3. Tricks for finding images of projections and singular points
To illustrate the method in the simplest case, let us suppose that $V$ has exactly one singular point. For almost all primes $p$, the reduction of $V$ mod $p$ will have the same property, and we can find the point mod $p$ by enumerating the ${\mathbb{F}}_{p}$-points of the reduction and checking them for singularity (deciding whether a point is singular is a simple computation if we know the dimension of $V$). By looking for rational numbers that reduce mod the various $p$ to the coordinates of these points, we guess the coordinates of the point over ${\mathbb{Q}}$. Again, this is not a proof, but we may easily check whether our candidate is actually on $V$ and, if so, whether it is singular.
@@
If the points are not rational, then of course for some primes we will only be able to find their reductions by passing to an extension of ${\mathbb{F}}_{p}$, which is undesirable. Fortunately in this work it was rarely necessary to find singular points not defined over a quadratic field. This case can be recognized by the singular points appearing on the reductions modulo half of the primes. There are a few more complicated cases with singular points defined over different fields; however, the quadratic fields of definition are easily recognized and so one can guess a formula for the number of singular points over ${\mathbb{F}}_{p}$ in terms of quadratic residue symbols. By considering the reductions modulo primes for which exactly one of the symbols is $1$ one can find the points.
@@ 3. Verification of modularity
Let us suppose that we have found a Weierstrass equation for the general fibre of an elliptic fibration. It may be defined by polynomials with unreasonably large coefficients, causing future computations to be slow. This situation can often be remedied as follows: find a minimal integral model and let $D(t)$ be the discriminant, ignoring any contribution at infinity so that this is a polynomial. Find an automorphism of $\mathbb{P}^{1}$ under which the numerator of the pullback of $D$ has smaller coefficients. (In Magma, this amounts to minimizing and reducing the hyperelliptic curve defined by $y^{2}=D(t)$.) Changing coordinates on the base by means of the same linear transformation often results in a much more usable equation, and twisting by a large square common factor of the coefficients is a further improvement.
@@ 4.1. Summary
In this paper we are mainly concerned with the cases where $p_{g}(H_{I,K})=1$. Up to Galois conjugacy, Hamahata [14] lists $14,6,4,2$ such surfaces over the real quadratic fields of discriminant $5,8,13,17$ respectively. (Such surfaces also exist over fields whose strict class number is not $1$, and over the fields of discriminant $29,37,41$ at level $1$, but these are not his concern.) We summarize his results on the Kodaira dimensions in the following table:
@@
Of the $7$ cases with $p_{g}=1$ where the surface was known not to be of general type, there are only $2$ where the level is not Galois-stable and thus the conjecture is not already known. As mentioned above, the surface $H_{{\mathfrak{p}}_{31},{\mathbb{Q}}(\sqrt{5})}$ has Kodaira dimension $1$, and we will show that $H_{2{\mathfrak{p}}_{7},{\mathbb{Q}}(\sqrt{2})}$, another surface left undecided by Hamahata, is of general type. These and other examples will be discussed in Section 5.
@@ 4.2. Level ${\mathfrak{p}}_{2}^{5}$ over ${\mathbb{Q}}(\sqrt{2})$
We do not know whether the generic fibre admits any ¯$\bar{\mathbb{Q}}(t)$-rational divisors of degree between 00 and $4$, but we suspect not.
@@ 4.3. Level $(3)$ over ${\mathbb{Q}}(\sqrt{13})$
We now present an example that, although still relatively simple, is more complicated than that of Section 4.2 and illustrates many of the methods of calculation used in this paper. Namely, we consider the isogeny class 2.2.13.1/9.1-a of elliptic curves of conductor $(3)$ over ${\mathbb{Q}}(\sqrt{13})$. Again, in this case the conductor is Galois-stable and so the Oda-Hamahata conjecture is already known. Nevertheless, the explicit calculation is of some interest.
@@ 4.5. Level $(2)$ over ${\mathbb{Q}}(\sqrt{17})$
In this section we consider a case where the level is Galois stable and thus the conjecture is known. Nevertheless there are some interesting questions to study.
@@ 4.6. Level $(3)$ over ${\mathbb{Q}}(\sqrt{2})$
Note that, as in the previous example, there is a prime $p>2$ dividing the Picard discriminant which is the isogeny degree between some curves in this isogeny class but does not divide the isogeny degree between a curve and its Galois conjugate. Thus, in order to find the desired correspondence over ${\mathbb{Q}}$, we would at some point need to use a map of degree divisible by $5$. In fact the Picard lattice is isomorphic to that of a K3 surface with $4$ fibres of type $I_{5}$, a section of height $8/5=4-2(6/5)$, and a $5$-torsion section. The $5$-isogeny goes to a surface of Picard discriminant $200$; however, as before it seems impossible to find a genus-$1$ fibration on this surface defined over ${\mathbb{Q}}$ with multisection degree $5$, so we cannot find the correspondence with a Kummer surface over ${\mathbb{Q}}$.
@@ 4.7. Level ${\mathfrak{p}}_{17}$ over ${\mathbb{Q}}(\sqrt{2})$
The minimal desingularization of $S_{5}$ has the same Hodge diamond as a K3 surface, but it is not itself a K3 surface. This follows, for example, from Kodaira’s formula for the canonical bundle of a surface with a genus-$1$ fibration, combined with the existence of a point of $\mathbb{P}^{1}$ where the fibre is nonreduced on $S_{5}$ but is a smooth rational curve on the Jacobian. The situation is analogous to that of an Enriques surface, for which the Hodge diamond is the same as that of a rational elliptic surface.
@@ 4.8. Level ${\mathfrak{p}}_{31}$ over ${\mathbb{Q}}(\sqrt{5})$
For the rest of this section let $\alpha=\frac{1+\sqrt{5}}{2}$. We will verify that the fibration is isomorphic to one on a K3 surface that is dominated by the Kummer surface $\mathcal{K}$ of $E\times E^{\sigma}$, where $E:y^{2}+xy+\alpha y=x^{3}+(\alpha+1)x^{2}+(\alpha-5)x+(3\alpha-5)$ is an elliptic curve of conductor ${\mathfrak{p}}_{31}$ (namely 2.2.5.1/31.1-a/5). Indeed, this curve is isomorphic to $y^{2}=x(x-1)(x-\frac{8\alpha+3}{16})$, so we may apply our general formulas to find a model for the Kummer surface over ${\mathbb{Q}}$ (taking care as usual to use the correct twist). There is a fibration with two $III^{*}$ fibres and no section, but since the $III^{*}$ fibres are not defined over ${\mathbb{Q}}$ this model cannot be directly compared with the one found as an image of the Hilbert modular surface.

### v2

@@ 1. Introduction
Let $E$ be an elliptic curve over a real quadratic field belonging to an isogeny class with one of the following LMFDB labels or their Galois conjugates: 2.2.5.1/31.1-a, 2.2.5.1/55.1-a, 2.2.5.1/80.1-a, 2.2.8.1/9.1-a, 2.2.8.1/17.1-a, 2.2.8.1/32.1-a, 2.2.13.1/9.1-a, 2.2.17.1/4.1-a, 2.2.17.1/8.3-a. Then there is a correspondence (not necessarily defined over ${\mathbb{Q}}$) between the Hilbert modular variety of level equal to the conductor of $E$ and the Kummer surface of $E\times E^{\sigma}$, inducing a map on $\operatorname{H}^{2}$ that is injective on the transcendental lattice of the Kummer surface.
@@
We now return to elliptic curves over $K$. Let $E$ be such a curve and let its conductor be $I$. To generalize the algebraic statement of modularity is fairly straightforward. One requires that the coefficients of the $L$-function of $E$ match the Hecke eigenvalues of a Hilbert modular form $f$. This is known for all elliptic curves over real quadratic fields and many over fields of higher degree: see [10, Theorems 1-2].
@@
[14, Section 0, p. 194] Let $E$ be an elliptic curve over a real quadratic field $K$ with conductor $I$. Then there is a correspondence between $H_{I,K}$ and $K_{E}=(E\times E^{\sigma})/\pm 1$ that induces an injection from the primitive part of $\operatorname{H}^{2}(K_{E})$ to that of $\operatorname{H}^{2}(H_{I,K})$.
@@
Hamahata does not define the term “primitive”, but we understand the phrase “primitive part of $\operatorname{H}^{2}$” to refer to the transcendental lattice, namely the quotient of $\operatorname{H}^{2}$ by the subspace spanned by fundamental classes of curves. We also remind the reader that a correspondence $X\subset V\times W$ induces maps on cohomology [37, p. 3]. We have $\operatorname{H}^{*}(V\times W)\cong\operatorname{H}^{*}(V)\otimes\operatorname{H}^{*}(W)$ by the Künneth formula, and $\operatorname{H}^{*}(V)\cong\operatorname{H}^{*}(V)^{\vee}$ by Poincaré duality. Thus $\operatorname{H}^{*}(V\times W)\cong\operatorname{Hom}(\operatorname{H}^{*}(V),\operatorname{H}^{*}(W))$, and one sees easily that if $X\subset V\times W$ with $\dim X=\dim V=\dim W$ (the only case used here) the map preserves degrees. Our correspondences will not be explicit subvarieties of $V\times W$; rather, they are induced from pairs of rational maps $\phi:V\dashrightarrow T,\psi:W\dashrightarrow T$ as the closure of the locus of $(v,w)$ where $\phi(v),\psi(w)$ are both defined and are equal.
Remark 1.12.
@@
Hamahata’s conjecture is known under some special circumstances: in particular, if $E$ is a ${\mathbb{Q}}$-curve, i.e., an elliptic curve isogenous to all of its Galois conjugates. This is not clearly stated or proved in the literature, but is known to experts. It is a consequence of the modularity of $E$, which is proved in [27, Corollary 6.2], subject to Serre’s conjecture on modular $2\times 2$ Galois representations, which is now a theorem of Khare and Wintenberger [17]. More precisely, Ribet’s result is that $E$ is a quotient of $J_{1}(N)$; it follows that there is a dominant map $J_{1}(N)^{d}\to\prod_{i=1}^{d}E^{\sigma_{i}}$, where $d=[K:{\mathbb{Q}}]$. The domain of this map is related to $\operatorname{H}^{d}(X_{0}(N))$ by Hecke correspondences, and the equality of the levels follows from Serre’s formula for the level of the modular form, also proved by Khare and Wintenberger.
@@
Let ${\mathcal{O}}$ be an order in a totally real field and ${\mathcal{O}}^{+}$ the set of its totally nonnegative elements. Let ${\mathfrak{d}}_{\mathcal{O}}$ be the different of ${\mathcal{O}}$. A Hilbert modular form for ${\mathcal{O}}$ has a Fourier expansion in which the coefficients are indexed by elements of ${\mathcal{O}}^{+}\cap{\mathfrak{d}}_{\mathcal{O}}^{-1}$ and are invariant under multiplication by elements of ${\mathcal{O}}^{+}\cap{\mathcal{O}}^{*}$. In order to multiply two such expansions, one needs an algorithm for writing an element of ${\mathcal{O}}^{+}$ as a sum of two such elements in all possible ways. An efficient method for doing so will be described in [1].
@@
In particular, this theorem applies in all cases where $H_{I,K}$ is not of general type and $I\neq(1)$, except for $K={\mathbb{Q}}(\sqrt{13}),I=(2)$. In this case we do find a K3 surface that is in correspondence with the Hilbert modular variety (and in fact this has been known for some years, being [35, Theorem 9]) but it is quite difficult to establish an explicit correspondence with the desired Kummer surface. In two of the cases where $I=(1)$ the birational equivalence between the Hilbert modular surface and a K3 surface with canonical singularities is made explicit in [38, Appendix].
@@
Our method is to study the Baily-Borel compactification $H_{I,K}$ by means of computer algebra and exhibit the desired correspondence to the Kummer surface by explicit maps. In some cases the Kummer surface of $E\times E^{\sigma}$ is itself a quotient of $H_{I,K}$; in other cases this does not appear to be true and instead we find an auxiliary K3 surface $S$ that admits maps of finite degree from both $H_{I,K}$ and the Kummer surface of $E\times E^{\sigma}$. The methods are similar to those of [38, Appendix], except that we do not necessarily start with a variety birational to a K3 surface. In addition, the setup of [38] requires information about the ring of Hilbert modular forms that is not available here: in particular, that it can be generated by Borcherds products. Against this, we do not find equations for Hirzebruch-Zagier cycles [36, Def. VI.1.4] on the Hilbert modular surfaces.
@@
In other cases, the surface or a quotient has geometric genus $1$ but is of general type. It is known [23, Theorem 1] that if a surface $S$ is of geometric genus $1$ then there is a K3 surface $K_{S}$ and a correspondence between $S$ and $K_{S}$ identifying the transcendental part of $H^{2}_{{\mathrm{\acute{e}t}}}$ for the two surfaces. However, finding the K3 surface and correspondence explicitly is a difficult and interesting problem in general. In some such cases the quotient by some or all of the Atkin-Lehner involutions is a surface with $p_{g}=1$ and $\kappa<2$ and we can find the desired correspondence, but otherwise there is no obvious method to use. This problem will be studied in future work.
@@ 2.2. Projective embedding and cusps
The projection away from a singular point or line may be a map of degree greater than $1$. This is undesirable, because the image may no longer have the desired cohomology. When it happens, we apply the $2$-fold Veronese embedding and proceed with our projections as before.
@@ 2.3. Tricks for finding images of projections and singular points
To illustrate the method in the simplest case, let us suppose that $V$ has exactly one singular point. For almost all primes $p$, the reduction of $V$ mod $p$ will have the same property, and we can find the point mod $p$ by enumerating the ${\mathbb{F}}_{p}$-points of the reduction and checking them for singularity (deciding whether a point is singular is a simple computation if we know the dimension of $V$). By looking for rational numbers that reduce modulo the various $p$ to the coordinates of these points, we guess the coordinates of the point over ${\mathbb{Q}}$. Again, this is not a proof, but we may easily check whether our candidate is actually on $V$ and, if so, whether it is singular.
@@
If the points are not rational, then for some primes we will only be able to find their reductions by passing to an extension of ${\mathbb{F}}_{p}$, which is undesirable. Fortunately in this work it was rarely necessary to find singular points not defined over a quadratic field. This case can be recognized by the singular points appearing on the reductions modulo half of the primes. There are a few more complicated cases with singular points defined over different fields; however, the quadratic fields of definition are easily recognized and so one can guess a formula for the number of singular points over ${\mathbb{F}}_{p}$ in terms of quadratic residue symbols. By considering the reductions modulo primes for which exactly one of the symbols is $1$ one can find the points.
@@ 3. Verification of modularity
Let us suppose that we have found a Weierstrass equation for the general fibre of an elliptic fibration. It may be defined by polynomials with unreasonably large coefficients, causing future computations to be slow. This situation can often be remedied as follows: find a minimal integral model and let $D(t)=\prod_{i}t-\alpha_{i}$, where the $\alpha_{i}$ are the coordinates of singular fibres excluding $\infty$. Find an automorphism of $\mathbb{P}^{1}$ under which the numerator of the pullback of $D$ has smaller coefficients. (In Magma, this amounts to minimizing and reducing the hyperelliptic curve defined by $y^{2}=D(t)$.) Changing coordinates on the base by means of the same linear transformation often results in a much more usable equation, and twisting by a large square common factor of the coefficients is a further improvement.
@@ 4.1. Summary
In this paper we are mainly concerned with the cases where $p_{g}(H_{I,K})=1$. Up to Galois conjugacy, Hamahata [14] lists $14,6,4,2$ such surfaces over the real quadratic fields of discriminant $5,8,13,17$ respectively. (Such surfaces also exist over fields whose strict class number is not $1$, and over the fields of discriminant $29,37,41$ at level $1$, but these are not his concern.) We summarize his results on the Kodaira dimensions.
Table 1. Kodaira dimensions of Hilbert modular surfaces with geometric genus $1$ over real quadratic fields with narrow class number $1$.
@@
Of the $7$ cases with $p_{g}=1$ where the surface was known not to be of general type, there are only $2$ that do not correspond to ${\mathbb{Q}}$-curves, so that the conjecture is not already known. As mentioned above, the surface $H_{{\mathfrak{p}}_{31},{\mathbb{Q}}(\sqrt{5})}$ has Kodaira dimension $1$, and we will show that $H_{2{\mathfrak{p}}_{7},{\mathbb{Q}}(\sqrt{2})}$, another surface left undecided by Hamahata, is of general type. These and other examples will be discussed in Section 5.
@@ 4.2. Level ${\mathfrak{p}}_{2}^{5}$ over ${\mathbb{Q}}(\sqrt{2})$
The generic fibre has no ${\mathbb{Q}}(t)$-rational divisors of degree between 00 and $4$, because there are fibres with no points over any quadratic extension of ${\mathbb{Q}}_{3}$. We do not know whether the generic fibre admits any ¯$\bar{\mathbb{Q}}(t)$-rational divisors of degree between 00 and $4$, but we suspect not.
@@ 4.3. Level $(3)$ over ${\mathbb{Q}}(\sqrt{13})$
We now present an example that, although still relatively simple, is more complicated than that of Section 4.2 and illustrates many of the methods of calculation used in this paper. Namely, we consider the isogeny class 2.2.13.1/9.1-a of elliptic curves of conductor $(3)$ over ${\mathbb{Q}}(\sqrt{13})$. Again, we have a ${\mathbb{Q}}$-curve and so the Oda-Hamahata conjecture is already known. Nevertheless, the explicit calculation is of some interest.
@@ 4.5. Level $(2)$ over ${\mathbb{Q}}(\sqrt{17})$
In this section we consider a ${\mathbb{Q}}$-curve, for which the conjecture is already known. Nevertheless there are some interesting questions to study.
@@ 4.6. Level $(3)$ over ${\mathbb{Q}}(\sqrt{2})$
Note that, as in the previous example, there is a prime $p>2$ dividing the Picard discriminant which is the isogeny degree between some curves in this isogeny class but does not divide the isogeny degree between a curve and its Galois conjugate. Thus, in order to find the desired correspondence over ${\mathbb{Q}}$, we would at some point need to use a map of degree divisible by $5$. In fact the Picard lattice is isomorphic to that of a K3 surface with $4$ fibres of type $I_{5}$, a section of height $8/5=4-2(6/5)$, and a $5$-torsion section. The $5$-isogeny goes to a surface of Picard discriminant $200$; however, as in Remark 4.30 it seems impossible to find a genus-$1$ fibration on this surface defined over ${\mathbb{Q}}$ with multisection degree $5$, so we cannot find the correspondence with a Kummer surface over ${\mathbb{Q}}$.
@@ 4.7. Level ${\mathfrak{p}}_{17}$ over ${\mathbb{Q}}(\sqrt{2})$
The minimal desingularization of $S_{5}$ has the same Hodge diamond as a K3 surface, but it is not itself a K3 surface. This follows from Kodaira’s formula for the canonical bundle of a surface with a genus-$1$ fibration, combined with the existence of a point of $\mathbb{P}^{1}$ where the fibre is nonreduced on $S_{5}$ but is a smooth rational curve on the Jacobian. The situation is analogous to that of an Enriques surface, for which the Hodge diamond is the same as that of a rational elliptic surface.
@@ 4.8. Level ${\mathfrak{p}}_{31}$ over ${\mathbb{Q}}(\sqrt{5})$
For the rest of this section let $\alpha=\frac{1+\sqrt{5}}{2}$. We will verify that the fibration is isomorphic to one on a K3 surface that is dominated by the Kummer surface $\mathcal{K}$ of $E\times E^{\sigma}$, where $E:y^{2}+xy+\alpha y=x^{3}+(\alpha+1)x^{2}+(\alpha-5)x+(3\alpha-5)$ is an elliptic curve of conductor ${\mathfrak{p}}_{31}$ (namely 2.2.5.1/31.1-a/5). Indeed, this curve is isomorphic to $y^{2}=x(x-1)(x-\frac{8\alpha+3}{16})$, so we may apply our general formulas to find a model for the Kummer surface over ${\mathbb{Q}}$, taking care as usual to use the correct twist. There is a fibration with two $III^{*}$ fibres and no section, but since the $III^{*}$ fibres are not defined over ${\mathbb{Q}}$ this model cannot be directly compared with the one found as an image of the Hilbert modular surface.